OpenStudy (anonymous):

The senior classes at High School A and High School B planned separate trips to the state fair. The senior class at High School A rented and filled 6 vans and 6 buses with 192 students. High School B rented and filled 8 vans and 16 buses with 456 students. Each van and each bus carried the same number of students. Find the number of students in each van and in each bus

OpenStudy (anonymous):

3x + 9y = 75 8x + 5y = 67 Multiply 1st equation by 8, and 2nd equation by 3 24x + 72y = 600 24x + 15y = 201 this is as far as i could go...i hope this is somewhat helpful...